Mary is a travel agent. She intends to sell her customers a special round-trip airline ticket package. She is able to purchase the package from the airline carrier for 170 € each. The round-trip tickets will be sold for 200 € each and the airline intends to reimburse Mary for any unsold ticket packages. Fixed costs include 5,140 € in advertising costs. . For every 135 round-trip airline ticket packages sold, operating income will increase by ________.

34,000 €

27,000 €

4,050 €

22,950 €
Answer #1

operating income will increase by 4050.

sales amount 135*200 = 27000

purchase value 135*170= 22950

operating income = 27000-22950 =4050